# Layer-Dependent Quantum Anomalous Hall Effect in Rhombohedral Graphene

### v1

@@ Abstract
In particular, we discovered that the QAH state is suppressed as the strength of Hund’s coupling $J_{\text{H}}$ increases. As we have discussed, the QAH state in this system can be modeled using the sign of the mass term of the four flavors, for example (+,+,+,)(+,+,+,-)( + , + , + , - ) for $(K\uparrow,K\downarrow,K^{\prime}\uparrow,K^{\prime}\downarrow)$. It is noted that the amplitudes of mass term of these four flavors are generally unequal, leading to different spin polarizations in the two valleys, for example $|\langle\mathbf{S}_{K^{\prime}}\rangle|\gg|\langle\mathbf{S}_{K}\rangle|\approx 0$. However, the Hund’s coupling tends to favor a ferromagnetic order with similar spin densities at two valleys. While in the above example, the spin density for the $K$ valley is much smaller than that for the $K^{\prime}$ valley. Consequently, the Hund’s coupling suppresses the QAH state as found in our numerical calculations.
@@
Fig. 2(b)-2(e) shows band structures of several typical phases. In the case of LAF-$x$, the colors represent the expectation value of spin polarization $\langle s_{z}\rangle$. For all other cases with spin $U(1)_{s}$ symmetry, the colors denote the layer polarization $\langle\rho_{z}\rangle$. Explicitly, we see that the band gap for LAF-$x$ state is larger than that of LAF-$z$ state. Furthermore, the layer polarization of the QAH state is consistent with with our analysis, where one flavor exhibits polarization opposite to the other three. This also indicates the QAH state here has Chern number $\mathcal{C}=5$. For the SM phase, it is originated from the warping effect of the band structure with one electron pocket. And depending on the parameters, it also features either one annular hole pocket or three smaller hole pockets.
@@
For thickness greater than five, the QAH state becomes SM and is no longer insulating due to the strong warping effect as shown in Fig. 3. In tetralayer graphene, the QAH state has a smaller gap compared to that in pentalayer for $\Delta<0$. Meanwhile, for $\Delta>0$, even though the QAH state now becomes SM, the indirect gap is much smaller in tetralyer. This is consistent with the experiment observations, where the anomalous Hall effect was found at zero magnetic filed, but only became quantized in the presence of a small magnetic field Sha et al. (2023). Remarkably, the critical field $\Delta_{c}$ for the LAF-$x$ state to SM transition is quantitatively consistent with experiments for various thicknessHan et al. (2023a); Liu et al. (2023); Han et al. (2023b); Sha et al. (2023); Zhou et al. (2023b). While for the second critical value of transition from SM to LP insulator, the disagreement may arise due to the fact that we may not simply relate $\Delta$ to the external field $\mathcal{D}$ without considering the screening effect Koshino (2010).
@@
Sandwich configuration. Finally, we consider the phase diagram of the sandwich configuration, namely TMD-graphene-TMD. Here the proximity Ising SOC would appear both in the top and bottom graphene layer with strength $\lambda^{t}$ and $\lambda^{b}$, respectively. The Ising SOC in TMD is odd under the $C_{2z}$ rotation. Thus, with two TMD materials neighboring the top and bottom layers, there are two configurations. One is the aligned case, where $\lambda^{t}$ and $\lambda^{b}$ have the same sign. The other is the anti-aligned case, with one TMD being rotated $180^{\circ}$ relative to the other, and $\text{sgn}(\lambda^{t})=-\text{sgn}(\lambda^{b})$. For simplicity, we set $|\lambda^{t}|=|\lambda^{b}|\equiv\lambda=1$ meV, and the Hamiltonian for these two cases is described by $\lambda\tau_{z}s_{z}$ and $\lambda\tau_{z}s_{z}\rho_{z}$, respectively. Fig. 4 shows the numerical phase diagram. For both cases, the phase diagram is symmetric respect to the sign of $\Delta$. For the aligned case, since Ising SOC acts cooperatively on the two surfaces, the QAH state is suppressed. While for the anti-aligned case, compared to the one-side proximity case, the only similar mass term still remains which enlarges the QAH state in the phase diagram.
Discussion. Our study provides a starting point for further theoretical and experimental investigation of correlated insulating phases in RMG with SOC. It also shows that in presence of Ising SOC, the system exhibits in-plane LAF order, which breaks spin $U(1)_{s}$ symmetry. Therefore, the system should have gapless collective excitation, namely the magnon. At the finite temperature, we would expect the Berezinskii–Kosterlitz–Thouless phase transition. Without SOC, the LAF state breaks the $SU(2)_{s}$ symmetry instead. It would be interesting to find the signature of the gapless magnon in future experiments, for instance through the spin transport experiment. Understanding the fluctuation effect to the Hartree-Fock mean field phase diagram is left for future work.
We also notice, experimentally, a $\mathcal{C}=3$ were stabilized by magnetic filed near the metal-insulator transition in pentalayer graphene Han et al. (2023a). Such a state can not be simply obtained from current mean field band theory. One possibility is the Hall crystal which spontaneously breaks the translation symmetry as studied in Ref. Dong et al. (2023), other may be from Landau level. We also leave all those to the future.
@@ Acknowledgements.
Acknowledgment. We acknowledge Guorui Chen for helpful discussions. This work is supported by the National Key Research Program of China under Grant No. 2019YFA0308404, the Natural Science Foundation of China through Grants No. 12350404 and No. 12174066, the Innovation Program for Quantum Science and Technology through Grant No. 2021ZD0302600, Science and Technology Commission of Shanghai Municipality under Grant No. 23JC1400600, and Shanghai Municipal Science and Technology Major Project under Grant No. 2019SHZDZX01.

### v2

@@ Abstract
In particular, we discovered that the QAH state is suppressed as the strength of Hund’s coupling $J_{\text{H}}$ increases. As we have discussed, the QAH state in this system can be modeled using the sign of the mass term of the four flavors, for example (+,+,+,)(+,+,+,-)( + , + , + , - ) for $(K\uparrow,K\downarrow,K^{\prime}\uparrow,K^{\prime}\downarrow)$. It is noted that the amplitudes of mass term of these four flavors are generally unequal, leading to different spin polarizations in the two valleys, for example $|\langle\mathbf{S}_{K^{\prime}}\rangle|\gg|\langle\mathbf{S}_{K}\rangle|\approx 0$. However, the Hund’s coupling tends to favor a ferromagnetic order with similar spin densities at two valleys. In the above example, the spin density for the $K$ valley is much smaller than that for the $K^{\prime}$ valley. Consequently, the Hund’s coupling suppresses the QAH state as found in our numerical calculations.
@@
Figs. 2(b)-2(e) show band structures of several typical phases. In the case of LAF-$x$, the colors represent the expectation value of spin polarization $\langle s_{x}\rangle$. For all other cases with spin $U(1)_{s}$ symmetry, the colors denote the layer polarization $\langle\varrho_{z}\rangle$. Explicitly, we see that the band gap for LAF-$x$ state is larger than that of LAF-$z$ state. Furthermore, the layer polarization of the QAH state is consistent with our analysis, where one flavor exhibits polarization opposite to the other three. This also indicates that the QAH state here has the Chern number $\mathcal{C}=5$. For the SM phase, it originated from the warping effect of the band structure with one electron pocket. Depending on the parameters, it also features either one annular hole pocket or three smaller hole pockets. Notably, we find that in the pentalayer system, even without SOC, the SM phase exhibits a non-vanishing anomalous Hall effect. Furthermore, based on orbital magnetization, a finite magnetic field with strength $B_{z}=1$ T can drive the system into a fully gapped QAH state, in agreement with experimental observations  [30, 64].
@@
For thickness greater than five, the QAH state becomes SM and is no longer insulating due to the strong warping effect as shown in Fig. 3. In tetralayer graphene, the QAH state has a smaller gap compared to that in pentalayer for $\Delta<0$. Meanwhile, for $\Delta>0$, even though the QAH state now becomes SM, the indirect gap is much smaller in tetralayer. This explains well the experimental observations, where the anomalous Hall effect was found at zero magnetic field, but only became quantized in the presence of a small magnetic field [34]. Remarkably, the critical field $\Delta_{c}$ for the LAF-$x$ state to SM transition is quantitatively consistent with experiments for various thicknesses[30, 31, 33, 34, 35]. While for the second critical value of transition from SM to LP insulator, the disagreement may arise due to the fact that we may not simply relate $\Delta$ to the external field $\mathcal{D}$ without considering the screening effect [68].
Sandwich configuration. Finally, we consider the phase diagram of the sandwich configuration, namely TMD-graphene-TMD. Here the proximity Ising SOC would appear in both the top and bottom graphene layers with strengths $\lambda^{t}$ and $\lambda^{b}$, respectively. The Ising SOC in TMD is odd under $C_{2z}$ rotation. Thus, with two TMD materials neighboring the top and bottom layers, there are two configurations. One is the aligned case, where $\lambda^{t}$ and $\lambda^{b}$ have the same sign. The other is the anti-aligned case[69], with one TMD being rotated $180^{\circ}$ relative to the other, and $\text{sgn}(\lambda^{t})=-\text{sgn}(\lambda^{b})$. For simplicity, we set $|\lambda^{t}|=|\lambda^{b}|\equiv\lambda=1$ meV, and the Hamiltonian for these two cases is described by $\lambda\tau_{z}s_{z}$ and $\lambda\tau_{z}s_{z}\varrho_{z}$, respectively. Fig. 4 shows the numerical phase diagram for pentalayer graphene. For both cases, the phase diagram is symmetric with respect to the sign of $\Delta$. For the aligned case, since Ising SOC acts cooperatively on the two surfaces, the QAH state is suppressed. Unlike the one-side proximity case, which induces the term $\lambda\tau_{z}s_{z}(\varrho_{0}-\varrho_{z})/2$, the anti-aligned case is described by $\lambda\tau_{z}s_{z}\varrho_{z}$. As a result, only the mass term from the one-side proximity case remains in the anti-aligned configuration, which enlarges the QAH state in the phase diagram. Therefore, such engineering of Ising SOC in RMG will stabilize the QAH effect with a full gap.
@@
Discussion. Our study provides a starting point for further theoretical and experimental investigation of correlated insulating phases in RMG with SOC. It also shows that in the presence of Ising SOC, the system exhibits an in-plane LAF order, which breaks the spin $U(1)_{s}$ symmetry. Therefore, the system should have gapless collective excitations, namely magnons. At finite temperature, the Berezinskii-Kosterlitz-Thouless phase transition is expected. Without SOC, the LAF state breaks the spin $SU(2)_{s}$ symmetry. It would be interesting to find the signature of the gapless magnon in future studies, for instance through spin transport. We notice, experimentally, that a $\mathcal{C}=3$ state was stabilized by a magnetic field near the metal-insulator transition in pentalayer graphene [30]. Such a state cannot be simply obtained from current mean field theory. One possibility is the Hall crystal, which spontaneously breaks the translation symmetry [70, 71, 72, 73, 74, 75, 76, 77, 78], another may be the Landau level. We leave these to future studies.
The interplay between interaction and SOC in proximitized RMG provides a versatile production line for the QAH effect with tunable Chern number. Remarkably, the rich choice of QAH states with different layer numbers, when proximitized to superconductivity, is promising to realize the exotic physics of chiral Majorana edge fermions and chiral Andreev edge states [79, 80, 81, 82, 83, 84, 85, 86].
@@ Acknowledgements.
Acknowledgment. We acknowledge Guorui Chen and Yuanbo Zhang for helpful discussions. This work is supported by the Natural Science Foundation of China through Grants No. 12350404 and No. 12174066, the Innovation Program for Quantum Science and Technology through Grant No. 2021ZD0302600, the Science and Technology Commission of Shanghai Municipality under Grants No. 23JC1400600, No. 24LZ1400100, and No. 2019SHZDZX01.